OpenStudy (happpily):

what value completes the square for the expression x^2+14x ?

OpenStudy (anonymous):

x why

OpenStudy (happpily):

what?

OpenStudy (anonymous):

(x+y)^2=x^2+2xy+y^2. if 2xy=14x, then y^2=___

OpenStudy (anonymous):

...????!!!

OpenStudy (happpily):

im so confused what.... ?!

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

U want the root?

OpenStudy (happpily):

honestly im not even 1005 what the questions asking..... ?

OpenStudy (happpily):

100%***

OpenStudy (anonymous):

i think it want u to find a number that make it s.t like this (x+7)^2=x^2+14x+49

OpenStudy (anonymous):

but i'm not sure

OpenStudy (happpily):

thanks for trying to helped out but i think ill just skip this question .-.